OpenStudy (anonymous):

how to take log of \[\beta ^{\alpha}(1+(x/\beta)^{\alpha})^{2}\]?

OpenStudy (anonymous):

\[\beta ^{\alpha}(1+(x/\beta)^{\alpha})^{2}\]

OpenStudy (anonymous):

\[ \log[\beta^\alpha (1+ (x/\beta)^\alpha)^2] = \log[\beta^\alpha] + \log[(1+(x/\beta)^\alpha)^2] \] \[ = \alpha\log[\beta] + 2\log[1+(x/\beta)^\alpha]\]